# Ego-to-Exo: Interfacing Third Person Visuals from Egocentric Views in Real-time for Improved ROV Teleoperation

### v1

@@ V Improved Underwater ROV Teleoperation
Multiple augmented viewpoints. We validate the utility of our proposed Ego-to-Exo teleoperation interface through further experiments on underwater cave exploration data. Our expedition in cave segments at Devil’s Springs, Florida reveals that when ROVs move slowly against strong currents, extending the viewpoint distance of the generated exocentric views can significantly improve teleoperation. This is achieved by tuning the queue parameters $r$, $c$, and $n$ in the proposed TeleOp interface. We consistently find that exocentric views are more informative, especially for about $5$-$10$ seconds preceding the ROV position during navigation. The multiple preceding views offered by our interface are particularly useful for mapping large structures such as newly discovered cave segments or shipwrecks. As Fig. 6 shows, the past ROV trajectories provide more spatial context in the augmented map, enabling operators to control the ROV efficiently around complex underwater structures.
Figure 6: A demonstration of our adjustable EOB viewpoint feature is shown. Teleoperators can slide across the distance ($f$) and find the best exocentric view, which is $f=200$ in this example.
Efficient teleoperation in complex missions. We also conducted trials inside an underwater cave system at Orange Grove, Florida, and a grotto system at Hudson, Florida with a total of five ROV operators. A BlueROV2 with a regular console was sent down for up to $300$-feet deep penetration inside the caves. When only egocentric views were available, the operators reported that maneuvering the robot by following the caveline was quite challenging because little/no ambient light penetrates inside underwater caves. Despite using powerful lights, problems such as moving shadows and scattered waves create significant blind spots. Consequently, tracking and following the caveline or any other navigation markers [46] without any peripheral view is extremely disorienting to the operator. In some cases, we observe that the cavelines get blended with the texture and features of cave walls in noisy conditions; see Fig. 7. In such scenarios, shifting the viewpoint to exocentric views with more global information allows easier identification of cavelines against the surrounding and overhead cave walls.
As Fig. 7 illustrates, our proposed Ego-to-Exo teleoperation interface generates exocentric views mimicking EOB views, which is helpful to operators for more efficient ROV teleoperation. Additionally, the augmented 3D map displays the robot’s pose, allowing much safer maneuvering of the vehicle to its desired orientation. Post-operation feedback from our ROV operators suggests that the exocentric EOB views are more useful for safe ROV maneuvers. They reported much better teleoperation experience as well as a significantly less cognitive load in conducting complex tasks such as object following and complex structure mapping.
@@ VI Conclusion & Future Work
This work presents an AR-based framework to synthesize exocentric camera views from egocentric feed in real-time for improved underwater ROV teleoperation. A pose geometry-based closed-form solution is formulated for the proposed Ego-to-Exo problem and then integrated into a monocular SLAM backbone. The end-to-end pipeline only requires a sequence of past egocentric views to generate an exocentric view with the accurate ROV model projected onto it. The geometric proof-of-concept is validated by ground plane estimation and reprojection error analyses in a series of 2D indoor navigation experiments. Subsequent field experiments are conducted for various challenging ROV teleoperation scenarios inside underwater caves at various locations. These experiments demonstrate that the proposed framework: (i) offers more informative peripheral views, (ii) provides better situational awareness, and (iii) facilitates an interactive ROV teleoperation experience. We are currently exploring more comprehensive multi-sensor fusion-based underwater SLAM backbone such as the SVIn2 [50] for more accurate and robust estimation. We further plan to develop and integrate more interactive features that would serve as a simulation platform for ROV teleoperation and navigation research.

### v2

@@ V Improved Underwater ROV Teleoperation
Multiple augmented viewpoints. We validate the utility of our proposed Ego-to-Exo teleoperation interface through further experiments on underwater cave exploration data. Our expedition in cave segments at Devil’s Springs, Florida reveals that when ROVs move slowly against strong currents, extending the exocentric viewpoint distance can significantly improve teleoperation. This is achieved by tuning the queue parameters $r$, $c$, and $n$ in the proposed TeleOp interface. We consistently find that exocentric views are more informative, especially for about $5$-$10$ seconds preceding the ROV position during navigation. The multiple preceding views offered by our interface are particularly useful for mapping large structures such as newly discovered cave segments or shipwrecks. As Fig. 6 shows, the synthesized viewpoints provide more spatial context, enabling operators to control the ROV efficiently around complex underwater structures.
Figure 6: A demonstration of our adjustable EOB viewpoint feature is shown. The $f$ numbers indicate the EOB distance from current to the reference frame used for the Ego-to-Exo view generation. Teleoperators can slide across the distance ($f$) and find the best exocentric view, which is $f=200$ in this example.
Efficient teleoperation in complex missions. We also conducted trials inside an underwater cave system at Orange Grove, Florida, and a grotto system at Hudson, Florida. We observe that maneuvering the robot by following the caveline with egocentric views is challenging because little/no ambient light penetrates inside underwater caves. Despite using powerful lights, problems such as moving shadows and scattered waves create significant blind spots. Consequently, tracking and following the caveline or any other navigation markers [49] without any peripheral view is extremely disorienting to the operator. In some cases, we observe that the cavelines get blended with the texture and features of cave walls in noisy conditions; see Fig. 7. In such scenarios, shifting the viewpoint to exocentric views allows easier identification of cavelines against the surrounding and overhead cave walls. Additionally, the augmented 3D map displays the robot’s pose, allowing much safer maneuvering of the vehicle to its desired orientation.
@@ VI Conclusion & Future Work
This work presents an AR-based framework to synthesize exocentric camera views from egocentric feed in real-time for improved underwater ROV teleoperation. A pose geometry-based closed-form solution is formulated for the proposed Ego-to-Exo problem and then integrated into a SLAM backbone. The end-to-end pipeline only requires a sequence of past egocentric views to generate an exocentric view with the accurate ROV model projected onto it. The proof-of-concept is validated by ground plane estimation and reprojection error analyses in a series of 2D indoor navigation experiments. Subsequent field experiments are conducted for various challenging ROV teleoperation scenarios inside underwater caves. A subjective study proves the advantage of the proposed TeleOp console over traditional systems and demonstrates that the framework: (i) offers more informative peripheral views, (ii) provides better situational awareness, and (iii) facilitates an interactive ROV teleoperation experience. We are currently exploring more comprehensive multi-sensor fusion-based underwater SLAM backbone such as the SVIn2 [54] for more accurate and robust estimation. We further plan to develop and integrate more interactive features that would serve as a simulation platform for ROV teleoperation and navigation research.